OpenStudy (anonymous):

i will medal There are 36,000 gallons of oil in 12 tank cars on the railroad track. Each tank car holds the same amount of oil. How many gallons of oil are in each tank car?

jhonyy9 (jhonyy9):

how you think it ?

jhonyy9 (jhonyy9):

write the rule of 3 simple and calculi the x value - ok. ? so than 36.000 gallons of oil in 12 tank cars x gallons of oil in 1 tank car ---------------------------------- cross multiplie and calculi the value of x x = (36000 *1)/12 = ?

OpenStudy (anonymous):

WHAT I HAVE NO IDEA WHAT THAT MEANS!!!???

OpenStudy (anonymous):

70?

jhonyy9 (jhonyy9):

36000/12= ? not will be 70 sure

jhonyy9 (jhonyy9):

so 36/12 = ? and there are again 3 zeros - yes ?